OpenStudy (anonymous):

solve by using quadratic form 2y squared-2y-1=0 the answer i got is way wrong, can anyone help me solve?

OpenStudy (anonymous):

You want to do this using the quadratic formula?

OpenStudy (anonymous):

thats what the book says to do

OpenStudy (anonymous):

Ok, first you need to compute b^2-4ac ay^2+by+c=0, so here a=2, b=-2 and c=-1

OpenStudy (anonymous):

want me to show you what i have and you can tell me where i went wrong?

OpenStudy (anonymous):

ok, that's fine too.

OpenStudy (mertsj):

\[2y^2-2y-1=0\] \[y=\frac{2\pm \sqrt{(-2)^2-4(2)(-1)}}{2(2)}=\frac{2\pm \sqrt{12}}{4}=\frac{2\pm2\sqrt{3}}{4}=\] \[\frac{1}{2}\pm \frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):